OpenStudy (anonymous):

partial differentiation question with complex numbers (i have very little experience, i have a few relevant books but have not started yet and im curious)

OpenStudy (anonymous):

does this form of chain rule work: \[\frac{dh( f(t) + g(t)i) }{dt} = h'( f(t) + g(t)i ).(f'(t) + g'(t) )\] i realise now its not really partial differentiation ( i dont think it is anyway)

OpenStudy (anonymous):

it works for some things i've tried it with, like \[ x^i\]

OpenStudy (experimentx):

looks like I am also illiterate on this matter.

OpenStudy (experimentx):

it could be a chain rule ... f(t) and g(t) being domain of h

OpenStudy (anonymous):

sorry, i meant : (x+i)^i

OpenStudy (experimentx):

http://www.physicsforums.com/showthread.php?t=53374 says that partial derivative for complex numbers work exactly like for real numbers.

OpenStudy (anonymous):

im pretty sure it works

OpenStudy (anonymous):

aw sweeet thats just what im looking for

OpenStudy (experimentx):

a complex number with complex power?? i haven't seen that.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%281%2Bi%29%5Ei there must be some way outta here

OpenStudy (experimentx):

let's change x+i into e's

OpenStudy (experimentx):

i guess that might solve our problem.

OpenStudy (anonymous):

aw they're pretty cool, when we did complex numbers for the first time last year i just spent lessons seeing how much i could figure out, so im not 100% on whether im allowed to do all this: \[a+bi = R(\cos\theta + isin\theta) = e^{\ln{R} + \theta i }\] \[\ln(a+bi) = \ln|a+bi| + Arg(a+bi)\] \[= \ln(\sqrt{a^2 + b^2}) + \arctan(b/a)\] \[(a + bi)^{c+di} = e^{(c+di)\ln({a+bi})}\] \[e^{(c+di)\ln(a+bi)} = e^{(c+di)(\ln(a^2 + b^2) + \arctan(b/a))}\] etc etc, it gets complicated quite quickly, so a chain rule would be very useful

OpenStudy (anonymous):

yeah thats the approach, using e to split it into f(x) + g(x)i

OpenStudy (experimentx):

i guess it does.

OpenStudy (phi):

@experimentX a complex number with complex power?? i haven't seen that. Then you might be interested in \( i^i \approx 0.2079\) http://www.wolframalpha.com/input/?i=I%5EI

OpenStudy (experimentx):

@phi truly interesting